# Gradient Extrapolation for Debiased Representation Learning

### v1

@@ Abstract
Machine learning classification models trained with empirical risk minimization (ERM) often inadvertently rely on spurious correlations. When absent in the test data, these unintended associations between non-target attributes and target labels lead to poor generalization. This paper addresses this problem from a model optimization perspective and proposes a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), designed to learn debiased representations in both known and unknown attribute training cases. GERNE uses two distinct batches with different amounts of spurious correlations to define the target gradient as the linear extrapolation of two gradients computed from each batch’s loss. It is demonstrated that the extrapolated gradient, if directed toward the gradient of the batch with fewer amount of spurious correlation, can guide the training process toward learning a debiased model. GERNE can serve as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases. The theoretical upper and lower bounds of the extrapolation factor are derived to ensure convergence. By adjusting this factor, GERNE can be adapted to maximize the Group-Balanced Accuracy (GBA) or the Worst-Group Accuracy. The proposed approach is validated on five vision and one NLP benchmarks, demonstrating competitive and often superior performance compared to state-of-the-art baseline methods.
@@ 1 Introduction
Deep learning models have demonstrated significant success in various classification tasks, but their performance is often compromised by datasets containing prevalent spurious correlations in the majority of samples [51, 29, 18, 13]. Spurious correlations refer to unintended associations between easy-to-learn non-target attributes and target labels, leading models based on Empirical Risk Minimization (ERM)— a widely used approach in classification tasks [44]— to rely on these correlations instead of the true, intrinsic features of the classes [12, 10, 40]. This occurs because the ERM objective optimizes for the average performance [44], which results in poor generalization when these spurious features are absent. For instance, in the Waterbirds classification task [45], where the goal is to classify a bird as either a waterbird or a landbird, the majority of waterbirds are associated with water backgrounds. In contrast, the majority of landbirds are associated with land backgrounds. A model trained with ERM might learn to classify the birds based on the background-water for waterbirds and land for landbirds-rather than focusing on the birds’ intrinsic characteristics. This reliance on the spurious feature allows the model to perform well on the majority training samples, where these correlations hold, but fails to generalize to test samples where these correlations are absent (e.g., waterbirds on land). Examples of Waterbirds images shown in Fig. 1$(a)$. Avoiding spurious correlations is crucial across various applications, including medical imaging [25, 37], finance [11], and climate modeling [17].
@@
This pervasive challenge has spurred extensive research into strategies for mitigating the negative effect of spurious correlations, particularly under varying levels of spurious attribute information availability. The authors of [49] provide a comprehensive review of the methods and research directions aimed at addressing this issue. In an ideal scenario, where attribute information is available in both the training and validation sets, methods can leverage this information to counteract spurious correlations [16, 39, 47]. When attribute information is available only in the validation set, methods either incorporate this set into the training process [18, 32, 42] or restrict its use to model selection and hyperparameter tuning [31, 28, 29, 27, 35]. Despite these efforts, existing methods still struggle to fully avoid learning spurious correlations, especially when the number of samples without spurious correlations is very limited in the training dataset, leading to poor generalization on the test data where these correlations are absent.
@@
In this paper, we adopt a different research approach, seeking to address the issue of spurious correlations from a model optimization perspective. We propose a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), to improve generalization and learn debiased representations. The contributions of this paper can be summarized as follows:
  • We propose GERNE, a novel and easy-to-implement debiasing method in classification tasks. The core idea is to sample two types of batches with varying amounts of spurious correlations and compute the two losses on these two batches. We linearly extrapolate the gradients of these two losses to obtain a target gradient. The target gradient, controlled by an extrapolation factor, is used to update the model’s parameters.
@@
  • The proposed gradient extrapolation approach is presented theoretically as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases.
  • The extrapolation factor’s theoretical upper and lower bounds are derived to ensure convergence, and its impact on performance is experimentally discussed.
  • We also establish a link between the extrapolation factor and both the Group-Balanced Accuracy (GBA) and Worst-Group Accuracy (WGA) metrics and generalize GERNE to the case with unknown attributes.
  • We highlight that in a biased dataset, overpresenting the minority groups in the sampled batches (compared to the majority) might be beneficial and can lead to SOTA results.
@@
  • We validate our approach on six benchmarks spanning both vision and NLP tasks, demonstrating superior performance compared to state-of-the-art methods.
@@
Figure 1: ($a$): Sample images from the waterbirds classification task. The majority of images with waterbirds have water background (e.g., $y=2,a=2$), and the majority of images with landbirds have land background (e.g., $y=1,a=1$). This correlation between bird class and background introduces spurious correlations in the dataset. $(b)$: $B_{b}$ shows a biased batch where the majority of images from class $y=1$ (first row of the batch) have an attribute $a=1$ (yellow color) and the majority of images from class $y=2$ (second row of the batch) have an attribute $a=2$ (light-blue color). $B_{lb}$ shows a more balanced attribute distribution (controlled by $c$, where $c=\frac{1}{2}$ in this example) than $B_{b}$ within each class. $B_{rs}$ shows a group-balanced distribution and refers to batch sampled using the Resampling method[16]. $B_{ext}$ simulates GERNE’s batch when $c\cdot(\beta+1)>1$ where the minority group in the dataset is represented as majority in this batch. $(c)$: A simplified 2D representation of gradient extrapolation (i.e. $\theta\in\mathbb{R}^{2}$). $\nabla_{\theta}\mathcal{L}_{b}$ represents the gradient computed on batch $B_{b}$, and training the model with this gradient is equivalent to training with ERM objective. $\nabla_{\theta}\mathcal{L}_{lb}$ represents the gradient computed on batch $B_{lb}$. $\nabla_{\theta}\mathcal{L}_{rs}$ is the gradient computed on batch $B_{rs}$ or equivalently, an extraplated gradient for $c\cdot(\beta+1)=1$. $\nabla_{\theta}\mathcal{L}_{ext}$ is our extrapolated gradient which can be controlled by the extrapolation factor $\beta$ depending the the level of bias in the data.
@@ 2 Related Work
Debiasing according to attributes annotations availability.
Numerous studies have leveraged attribute annotations to mitigate spurious correlations and learning debiased representation [39, 54, 3, 50, 52]. For instance, Group DRO [39] optimizes model performance on the worst-case group by minimizing worst-group error during training. While these methods are effective, obtaining attribute annotations for each sample can be extremely time-consuming and labor-intensive. Consequently, recent works have explored approaches that rely on limited attribute-labeled data to reduce the dependency on full annotations [18, 42, 32]. For example, DFR [18] enhances robustness by using a small, group-balanced validation set with attribute labels to retrain the final layer of a pre-trained model. For cases where attribute information is only available for model selection and hyperparameter tuning [6, 16, 28, 31, 53], usually an initial model is used to separate samples based on the alignment between the label and spurious attributes. Samples for which this model incurs relatively low loss are considered ”easy” examples, where we expect the alignment to hold and the samples to closely resemble the majority group. In contrast, samples with high loss are considered ”hard” examples, and these samples tend to resemble the minority group [48]. This process effectively creates ”easy” and ”hard” pseudo-attributes within each class, allowing debiasing methods that traditionally rely on attribute information to be applied. For example, JTT [28] first trains a standard ERM model and then trains a second model by upweighting the misclassified training examples (hard examples) detected by the first model. Finally, a more realistic and challenging scenario arises when attribute information is entirely unavailable [4, 43]—not accessible for training, model selection, or hyperparameter tuning—requiring models to generalize without explicit guidance on non-causal features [49].
@@
A prominent family of solutions to mitigate spurious correlations across the aforementioned scenarios of annotation availability involves data balancing techniques [19, 16, 46, 40, 36, 7, 21]. These methods are valued for their simplicity and adaptability, as they are typically faster to train and do not require additional hyperparameters. Resampling underrepresented groups to ensure a more balanced distribution of samples [16, 19] or modifying the loss function to adjust for imbalances [38] are common examples of these techniques.
In this work, we demonstrate in Sec. 5.3 that although balancing techniques are effective, their performance is constrained in the presence of spurious correlations. In contrast, our proposed debiasing approach mitigates the negative effects of spurious correlations by guiding the learning process in a debiasing direction, proving to be more effective.
@@ 3 Problem Setup
We consider a regular multi-class classification problem with $K$ classes and $A$ spurious attributes. Each input sample $x_{i}\in\mathcal{X}=\{x_{i}\mid i=1,\dots,N\}$ is associated with a class label $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$ and an attribute $a_{i}\in\mathcal{A}=\{1,\dots,A\}$, where $N$ is the total number of input samples in the dataset. We define a group $\mathcal{X}_{y,a}$ for $(y,a)\in\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ as the set of input samples $x_{i}$ with a class label $y$ and an attribute $a$, resulting in $|\mathcal{G}|=K\cdot A$ groups. We define $\mathcal{X}_{y}=\cup_{a\in\mathcal{A}}\mathcal{X}_{y,a}$ as the set of input samples $x_{i}$ with a class label $y$. We assume that all groups are non-empty (i.e. $\forall(y,a)\in\mathcal{G},\mathcal{X}_{y,a}\neq\emptyset$).
Our goal is to learn the intrinsic features that define the labels rather than spurious features present in a biased dataset. This ensures robust generalization when spurious correlations are absent in the test distribution. Following [39], we aim to learn a function parameterized by a neural network $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to minimize the risk for the worst-case group:
$f^{*}=\arg\min_{f}\max_{g\in\mathcal{G}}\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right].$ (1)
@@ 4 The Proposed Method: GERNE
The core idea of GERNE is to sample two batches with different amounts of spurious correlations, hereafter named the biased batch $\boldsymbol{B_{b}}$ and the less biased batch $\boldsymbol{B_{lb}}$ (Fig. 1$(b)$). Let $\mathcal{L}_{b},\mathcal{L}_{lb}$ be the losses calculated on $\boldsymbol{B_{b}}$ and $\boldsymbol{B_{lb}}$, respectively. We assume that utilizing the extrapolation of the gradients of these two losses towards the gradient of $\mathcal{L}_{lb}$ guides the model toward debiasing as illustrated in Fig. 1$(c)$. We first present our method GERNE for training with known attributes and then generalize GERNE to the unknown attribute case.
@@ 4.1 GERNE for the Known Attributes Case
In the following, we denote the joint distribution of $(y,a)$ in a sampled batch from the dataset as $p(y,a)$. We define two types of batches with different conditional attribute distributions $p(a|y)$: the biased and the less biased batches. Then, we define our target loss as the linear extrapolation of the two losses computed on the biased and less biased batches. A simplified illustration of our approach is shown in Fig. 1. Finally, we derive the link between the extrapolation factor and the risk of the worst-case group in Eq. 1, and we theoretically define the upper and lower bounds of this factor.
@@ 4.1.1 Sampling the biased and the less biased batches
The biased batch and less biased batches are sampled to satisfy the following two conditions:
@@
    Uniform sampling from classes, i.e., $\forall y\in\mathcal{Y},p(y)=\frac{1}{K}$
@@
Here $|\mathcal{X}_{y,a}|,|\mathcal{X}_{y}|$ denote the cardinality of $\mathcal{X}_{y,a},\mathcal{X}_{y}$, respectively. Note that to sample a biased batch, no access to the attributes is required, and uniformly sampling from $\mathcal{X}_{y}$ for each label $y$ satisfies Eq. 2. While the less biased batch ($\boldsymbol{B_{lb}}$) is sampled with a conditional attribute distribution, denoted as $p_{lb}(a|y)$, which is defined to be less biased compared to the biased batch. That is: $\forall(y,a)\in\mathcal{G}$,
@@
Here $\mathcal{L}_{lb}$ quantifies the loss when the representation of the spurious features is partially reduced. Choosing
@@ 4.1.2 Gradient extrapolation
$\mathcal{L}_{ext}=\mathcal{L}_{lb}+\beta(\mathcal{L}_{lb}-\mathcal{L}_{b}).$ (5)
@@
Given the set of parameters $\theta$ of our model $f$, the gradient vector of the loss $\mathcal{L}_{ext}$ with respect to $\theta$ can be derived from Eq. 5:
$\nabla_{\theta}\mathcal{L}_{ext}=\nabla_{\theta}\mathcal{L}_{lb}+\beta\left(\nabla_{\theta}\mathcal{L}_{lb}-\nabla_{\theta}\mathcal{L}_{b}\right).$ (7)
Our target gradient vector $\nabla_{\theta}\mathcal{L}_{ext}$ in Eq. 7 is the linear extrapolation of the two gradient vectors $\nabla_{\theta}\mathcal{L}_{lb}$ and $\nabla_{\theta}\mathcal{L}_{b}$, and accordingly, we call $\beta$ the extrapolation factor. Since the less biased batch has a less skewed conditional attribute distribution in comparison to the biased batch according to Eq. 3, the extrapolation of their gradients and toward the less biased gradient forms a new gradient ($\mathcal{L}_{ext}$) that leads to learning even more debiasing representation for some values of the extrapolation factor $\beta>0$. The concept of extraploation is shown in Fig. 1$(c)$.
4.1.3 GERNE as general framework for debiasing

### v2

@@ Abstract
Machine learning classification models trained with empirical risk minimization (ERM) often inadvertently rely on spurious correlations. When absent in the test data, these unintended associations between non-target attributes and target labels lead to poor generalization. This paper addresses this problem from a model optimization perspective and proposes a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), designed to learn debiased representations in both known and unknown attribute training cases. GERNE uses two distinct batches with different amounts of spurious correlations and defines the target gradient as a linear extrapolation of the gradients computed from each batch’s loss. Our analysis shows that when the extrapolated gradient points toward the batch gradient with fewer spurious correlations, it effectively guides training toward learning a debiased model. GERNE serves as a general framework for debiasing, encompassing ERM and Resampling methods as special cases. We derive the theoretical upper and lower bounds of the extrapolation factor employed by GERNE. By tuning this factor, GERNE can adapt to maximize either Group-Balanced Accuracy (GBA) or Worst-Group Accuracy (WGA). We validate GERNE on five vision and one NLP benchmarks, demonstrating competitive and often superior performance compared to state-of-the-art baselines. The project page is available at: https://gerne-debias.github.io/.
@@ 1 Introduction
Deep learning models have demonstrated significant success in various classification tasks, but their performance is often compromised by datasets containing prevalent spurious correlations in the majority of samples  [52, 29, 18, 13]. Spurious correlations refer to unintended associations between easy-to-learn, non-target attributes and target labels. These correlations often cause models trained with Empirical Risk Minimization (ERM) [45] to rely on these correlations instead of the true, intrinsic features of the classes  [12, 10, 40]. This occurs because the ERM objective optimizes for the average performance  [45], thereby biasing the model toward the easy-to-learn features that are predictive for the majority of training samples. As a result, ERM-trained models often exhibit poor generalization when these spurious features are absent in the test data. For instance, in the Waterbirds classification task  [46], where the goal is to classify a bird as either a waterbird or a landbird, the majority of waterbirds are associated with water backgrounds. In contrast, the majority of landbirds are associated with land backgrounds. A model trained with ERM might learn to classify the birds based on the background—water for waterbirds and land for landbirds—rather than focusing on the birds’ intrinsic characteristics. This reliance on the spurious feature allows the model to perform well on the majority of training samples, where these correlations hold, but fails to generalize to test samples where these correlations are absent (e.g., waterbirds on land). Examples of Waterbirds images shown in Fig. 1(a). Avoiding spurious correlations is crucial across various applications, including medical imaging  [25, 37], finance  [11], and climate modeling  [17].
@@
This pervasive challenge has spurred extensive research into strategies for mitigating the negative effects of spurious correlations, particularly under varying levels of attribute information availability. The authors of  [50] provide a comprehensive review of the methods and research directions aimed at addressing this issue. In an ideal scenario, where attribute information is available in both the training and validation sets, methods can leverage this information to counteract spurious correlations  [16, 39, 48]. When attribute information is available only in the validation set, methods either incorporate this set into the training process  [18, 32, 42] or restrict its use to model selection and hyperparameter tuning  [31, 28, 29, 27, 35]. Despite these efforts, existing methods still struggle to fully avoid learning spurious correlations, especially when the number of samples without spurious correlations is very limited in the training dataset, leading to poor generalization on the test data where these correlations are absent.
@@
In this paper, we adopt a different research approach, seeking to address the issue of spurious correlations from a model optimization perspective. We propose a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), to avoid reliance on spurious features and learn debiased representations. The core idea is to sample two types of batches with varying amounts of spurious correlations (Fig. 1(b)) and compute the two losses on these two batches. We assume that the difference between the gradients of these losses captures a debiasing direction. Therefore, we define our target gradient as the linear extrapolation of these two gradients toward the gradient of the batch with fewer amount of spurious correlations (Fig. 1(c)). The contributions of this paper can be summarized as follows:
@@
  • We propose GERNE as a general framework for debiasing, with methods such as ERM and Resampling being shown as special cases.
  • We derive the theoretical upper and lower bounds of the extrapolation factor and establish a direct connection between the extrapolation factor and the risk for the worst-case group. We show that tuning this factor within these bounds enables GERNE to adaptively optimize for either Group-Balanced Accuracy or Worst-Group Accuracy.
@@
  • We validate our approach on six benchmarks spanning both vision and NLP tasks, under both known and unknown attribute cases, demonstrating competitive and often superior performance compared to state-of-the-art methods—particularly in scenarios where samples without spurious correlations are scarce.
@@
Figure 1: 1(a): Sample images from the waterbirds classification task. Most landbird images appear with land backgrounds (i.e., $y=1,a=1$), while most waterbird images appear with water backgrounds (i.e., $y=2,a=2$). This correlation between bird class and background introduces spurious correlations in the dataset. 1(b): Visualization of batch construction. $B_{b}$ refers to a biased batch where the majority of images from class $y=1$ (top row) have attribute $a=1$ (yellow), and the majority of images from class $y=2$ (bottom row) have attribute $a=2$ (light-blue). $B_{lb}$ represents a less biased batch, with a more balanced attribute distribution within each class, controlled by $c$ (here $c=\frac{1}{2}$). $B_{rs}$ depicts a batch with group-balanced distribution and refers to the batch used in the Resampling method [16]. $B_{ext}$ simulates GERNE’s batch with $c\cdot(\beta+1)>1$, where the dataset’s minority groups appear as majorities in the batch. 1(c): A simplified 2D representation of gradient extrapolation where $\theta\in\mathbb{R}^{2}$. $\nabla_{\theta}\mathcal{L}_{b}$ is the gradient computed on $B_{b}$; training with this gradient is equivalent to training with the ERM objective. $\nabla_{\theta}\mathcal{L}_{lb}$ represents the gradient computed on $B_{lb}$. $\nabla_{\theta}\mathcal{L}_{rs}$ is the gradient computed on $B_{rs}$, which is equivalent, in expectation, to an extrapolated gradient with $c\cdot(\beta+1)=1$. Finally, $\nabla_{\theta}\mathcal{L}_{ext}$ is our extrapolated gradient, with the extrapolation factor $\beta$ modulating the degree of debiasing based on the strength of spurious correlations present in the dataset.
@@ 2 Related Work
Debiasing according to attribute annotations availability.
Numerous studies have leveraged attribute annotations to mitigate spurious correlations and learning debiased representation [39, 55, 3, 51, 53]. For instance, Group DRO [39] optimizes model performance on the worst-case group by directly minimizing worst-group loss during training. While effective, such methods rely on complete attribute annotations, which are often costly and labor-intensive to obtain. Consequently, recent works have explored approaches that reduce reliance on full annotations by using limited attribute information [18, 42, 32]. For example, DFR [18] enhances robustness by using a small, group-balanced validation set with attribute information to retrain the final layer of a pre-trained model. In cases where attribute information is only available for model selection and hyperparameter tuning [6, 16, 28, 31, 54], an initial or auxiliary ERM-trained model is often used to infer the attributes by partitioning the training data into majority and minority groups. Samples on which the model incurs relatively low loss (i.e., high-confidence predictions) are treated as “easy” examples—where spurious correlations are likely to hold— and these examples form the majority group. Conversely, high-loss samples are considered “hard” examples, and typically form the minority group where such correlations may not apply [49]. This process effectively creates “easy” and “hard” pseudo-attributes within each class, allowing debiasing methods that traditionally rely on attribute information to be applied. For example, JTT [28] first trains a standard ERM model and then trains a second model by upweighting the misclassified training examples detected by the first model. Finally, a more realistic and challenging scenario arises when attribute information is entirely unavailable [4, 43]—not accessible for training, model selection, or hyperparameter tuning—requiring models to generalize without explicit guidance on non-causal features [50].
@@
A prominent family of solutions to mitigate spurious correlations across the aforementioned scenarios of annotation availability involves data balancing techniques [19, 16, 47, 40, 36, 7, 21]. These methods are valued for their simplicity and adaptability, as they are typically faster to train and do not require additional hyperparameters. Resampling underrepresented groups to ensure a more balanced distribution of samples [16, 19] or modifying the loss function to adjust for imbalances [38] are common examples of these techniques. We demonstrate in Sec. 5.3 that although the balancing techniques are effective, their performance is constrained in the presence of spurious correlations. In contrast, our proposed debiasing approach mitigates the negative effects of spurious correlations by guiding the learning process in a debiasing direction, proving to be more effective.
@@ 3 Problem Setup
We consider a standard multi-class classification problem with $K$ classes and $A$ attributes. Each input sample $x_{i}\in\mathcal{X}=\{x_{j}\mid j=1,\dots,N\}$ is associated with a class label $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$ and an attribute $a_{i}\in\mathcal{A}=\{1,\dots,A\}$, where $N$ is the total number of samples in the dataset. We define a group $\mathcal{X}_{y,a}$ for $(y,a)\in\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ as the set of input samples $x_{i}$ with class label $y$ and attribute $a$, resulting in $|\mathcal{G}|=K\cdot A$ groups. For each class $y$, we denote by $\mathcal{X}_{y}=\bigcup_{a\in\mathcal{A}}\mathcal{X}_{y,a}$ the set of all samples with label $y$. We assume all groups are non-empty, i.e., $\forall(y,a)\in\mathcal{G},\mathcal{X}_{y,a}\neq\emptyset$, and denote the cardinality of any group $\mathcal{X}_{m}$ by $|\mathcal{X}_{m}|$.
Our goal is to learn the intrinsic features that define the labels, rather than spurious features present in a biased dataset, where spurious correlations are prevalent. This would ensure robust generalization when spurious correlations are absent in the test distribution. Following  [39], we aim to learn a function parameterized by a neural network $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to minimize the risk for the worst-case group:
$f^{*}=\arg\min_{f}\max_{g\in\mathcal{G}}\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right],$ (1)
@@ 4 The Proposed Method: GERNE
We build GERNE with the goal of mitigating the impact of spurious correlations. The core idea of GERNE is to sample two batches with different amounts of spurious correlations, hereafter named the biased batch $\boldsymbol{B_{b}}$ and the less biased batch $\boldsymbol{B_{lb}}$ (Fig. 1(b)). Let $\mathcal{L}_{b},\mathcal{L}_{lb}$ be the losses calculated on $\boldsymbol{B_{b}}$ and $\boldsymbol{B_{lb}}$, respectively. We assume that extrapolating the gradients of these two losses towards the gradient of $\mathcal{L}_{lb}$ guides the model toward debiasing as illustrated in Fig. 1(c). We first present GERNE for training with known attributes and then generalize GERNE to the unknown attribute case.
@@ 4.1 GERNE for the Known Attributes Case
In the following, we denote by $p(y,a)$ the joint distribution of class label $y$ and attribute $a$ in a sampled batch. During training, we construct two types of batches with different conditional attribute distributions $p(a|y)$: the biased and the less biased batches. Our method defines the target loss as a linear extrapolation between the losses computed on these two batches. A simplified illustration is shown in Fig. 1. Finally, we derive the link between the extrapolation factor and the risk for the worst-case group in Eq. 1, and theoretically define the upper and lower bounds of this factor.
@@ 4.1.1 Sampling the biased and the less biased batches
The biased batch and the less biased batches are sampled to satisfy the following two conditions:
@@
    Uniform sampling from classes, i.e., $\forall y\in\mathcal{Y},p(y)=\frac{1}{K}.$
@@
Note that to sample a biased batch, no access to the attributes is required, and uniformly sampling from $\mathcal{X}_{y}$ for each label $y$ satisfies Eq. 2. The less biased batch ($\boldsymbol{B_{lb}}$) is sampled with a conditional attribute distribution, denoted as $p_{lb}(a|y)$, which satisfies the following: $\forall(y,a)\in\mathcal{G}$:
@@
That is, $B_{lb}$ exhibits a more balanced group distribution than $B_{b}$, and $\mathcal{L}_{lb}$ quantifies the loss when spurious correlations are reduced in the sampled batch. Choosing
@@ 4.1.2 Gradient extrapolation
$\mathcal{L}_{ext}=\mathcal{L}_{lb}+\beta\cdot(\mathcal{L}_{lb}-\mathcal{L}_{b}),$ (5)
@@
Given the set of parameters $\theta$ of our model $f$, the gradient of $\mathcal{L}_{ext}$ with respect to $\theta$ can be derived from Eq. 5:
$\nabla_{\theta}\mathcal{L}_{ext}=\nabla_{\theta}\mathcal{L}_{lb}+\beta\cdot\left(\nabla_{\theta}\mathcal{L}_{lb}-\nabla_{\theta}\mathcal{L}_{b}\right).$ (7)
Our target gradient vector $\nabla_{\theta}\mathcal{L}_{ext}$ in Eq. 7 is a linear extrapolation of the two gradient vectors $\nabla_{\theta}\mathcal{L}_{lb}$ and $\nabla_{\theta}\mathcal{L}_{b}$, and accordingly, we refer to $\beta$ as the extrapolation factor. Because the less biased batch has a less skewed conditional attribute distribution compared to the biased batch (as shown in Eq. 3), extrapolating their gradients and toward the less biased gradient forms a new gradient ($\mathcal{L}_{ext}$) that leads to learning even more debiased representation for some values of the extrapolation factor $\beta>0$. A visual representation of extrapolation is shown in Fig. 1(c).
4.1.3 GERNE as a general framework for debiasing